A penny is dropped into a tank of water at the water’s surface. If falls to the bottom according to the relation below , where d is the depth of the water measured in metres and t is the time after the penny was dropped, measured in seconds. How deep is the water?



1
Expert's answer
2022-03-03T11:51:35-0500

The question is incomplete the relation equation is not provided. However, I found a similar question with the relation equation being d=3.5t2+35.

Question

A penny is dropped into a tank of water's surface. It falls to the bottom according to the relation d=3.5t^2+35, where d is the depth of the water measured in meters and t is the time after the penny was dropped, measured in seconds. How deep is the water?


Hence the depth of the water depends on the time the penny takes to reach the bottom of the tank


If t = 0 sec, then d=3.5 x 0+ 35 = 35 m

If t = 1 sec, then d=3.5 x 1+ 35 = 38.5 m

If t = 2 sec, then d=3.5 x 2+ 35 = 49 m

If t = 3 sec, then d=3.5 x 3+ 35 = 66.5 m

If t = 4 sec, then d=3.5 x 4+ 35 = 91 m

If t = 5 sec, then d=3.5 x 5+ 35 = 122.5 m

If t = 6 sec, then d=3.5 x 6+ 35 = 161 m

If t = 7 sec, then d=3.5 x 7+ 35 = 206.5 m

If t = 8 sec, then d=3.5 x 8+ 35 = 259 m

If t = 9 sec, then d=3.5 x 9+ 35 = 318.5 m

If t = 10 sec, then d=3.5 x 10+ 35 = 385 m [Answer]